### v3

# H2R: A Human-to-Robot Data Augmentation for Robot Pre-training from Videos

Abstract
Large-scale pre-training using videos has proven effective for robot learning. However, the models pre-trained on such data can be suboptimal for robot learning due to the significant visual gap between human hands and those of different robots. To remedy this, we propose H2R, a simple data augmentation technique that detects human hand keypoints, synthesizes robot motions in simulation, and composites rendered robots into egocentric videos. This process explicitly bridges the visual gap between human and robot embodiments during pre-training. We apply H2R to augment large-scale egocentric human video datasets such as Ego4D and SSv2, replacing human hands with simulated robotic arms to generate robot-centric training data. To verify the effectiveness of the augmentation pipeline, we introduce a CLIP-based image-text similarity metric that quantitatively evaluates the semantic fidelity of robot-rendered frames to the original human actions. We validate H2R across four simulation benchmarks—Robomimic, RLBench, PushT and CortexBench, and real-world manipulation tasks with UR5 and Dual-Arm Franka/UR5. H2R consistently improves downstream success rates, yielding gains of 1.3%–10.2% in simulation and 3.3%–23.3% in real-world tasks across various visual encoders and policy learning methods. We further demonstrate the potential of H2R in cross-embodiment generalization and its compatibility with VLA models. These results indicate that H2R improves the generalization ability of robotic policies by mitigating the visual discrepancies between human and robot domains. The project page is available in https://sites.google.com/view/h2r-robotics.
1 Introduction
Pre-training of generalizable robotic features for object manipulation and motion navigation constitutes a crucial objective within the realm of robotics. Inspired by the remarkable accomplishments of large scale pre-training in computer vision [he2022mae, nair2022r3m, zeng2024learningmanipulationpredictinginteraction, karamcheti2023voltron, datacomp] and natural language processing [devlin2018bert, yu2022pointbert, openai2024gpt4technicalreport, anil2023palm2technicalreport, li2024datacomplm], many efforts have been devoted to harnessing large-scale data to construct generalizable representations in the robotics field [embodimentcollaboration2024openxembodimentroboticlearning, brohan2023rt1roboticstransformerrealworld, brohan2023rt2visionlanguageactionmodelstransfer]. Nevertheless, when it comes to robot manipulation, the process of collecting demonstrations is labor-intensive and expensive [duan2023ar2, jang2022bczzeroshottaskgeneralization, brohan2023rt1roboticstransformerrealworld, khazatsky2024droidlargescaleinthewildrobot, fang2023rh20tcomprehensiveroboticdataset, gao2024efficientdatacollectionrobotic, lee2021pickandplacetacklingroboticstacking, herzog2023deeprlscalesorting, kalashnikov2021mtoptcontinuousmultitaskrobotic, duan2024manipulate]; meanwhile, there exist many large-scale egocentric video datasets showing how humans perform manipulation and navigation, which can potentially serve as a cheap alternative of demonstrations for the pre-training of generalizable visual features for robotics.
Recent works [nair2022r3m, xiao2022mvp, zeng2024learning] analyze such egocentric human video datasets such as Ego4D [grauman2022ego4d], SSv2 [goyal2017something], and Epic Kitchens [damen2018scaling] with the aim of gleaning prior knowledge about object manipulation and enabling the acquisition of general and robust feature representations. However, during the representation learning, the gap in visual representations between the human arm and the robotic arm remains largely unaddressed and can hinder the transferability of models trained on egocentric datasets to robotic systems. Specifically, when utilizing the robot expert data to fine-tune the pre-trained robotic representations for downstream robotic tasks, the model has to learn to bridge the visual gap between the first-person human hand and the robots in addition to acquiring nuanced task-specific skills demonstrated in the robot expert data. This would result in increased complexity during the fine-tuning process and suboptimal performance.
To mitigate this issue, we propose H2R (as shown in Figure LABEL:fig:pipeline), a simple data augmentation method that converts videos of Human hand operations into that of Robotic arm manipulation. H2R consists of two major procedures: the first part is to generate the robotic movements to imitate the human hand movements in a video, followed by the second stage that overlays the robotic movements onto the human hand’s movements in the video. Specifically, in the first part, we employ state-of-the-art 3D hand reconstruction model HaMeR [pavlakos2023reconstructinghands3dtransformers] to accurately detect the position and posture of the human hand in egocentric videos. Then, we simulate the same robot state in simulators to obtain the mask of robot actions. In the second stage, we use the Segment Anything Model [kirillov2023segment] to automatically separate human hand from background, and use the inpainting model LaMa [suvorov2021resolutionrobustlargemaskinpainting] to fill the removed hand mask. After that, we align the camera intrinsic parameters of the images detected in HaMeR with those in the simulator, and then achieve pixel-level matching between the robotic arm images in the simulators and the human hand images in the egocentric video. Finally, we overlay the robotic arm images captured by the simulator’s camera onto the areas where the human hands are removed. Through such a process, H2R explicitly reduces the gap between human and robot hands by creating realistic robotic arm movements that visually mimic human hand actions. It allows the model to learn the task-specific actions demonstrated by the human hand, but with robotic arm visual representations that are more suitable for robotic systems.
To evaluate the effectiveness of the H2R augmentation process, we introduce a CLIP-based semantic similarity metric that measures how well the rendered robot frames preserve the original action semantics. This provides a lightweight and scalable proxy to assess the alignment quality between input human videos and robot-augmented outputs.
To further verify the utility of the augmented datasets, we conduct downstream experiments comparing models pre-trained on original egocentric datasets with those trained on our released H2R-enhanced datasets. These comparisons are performed on both simulation and real-world robotic manipulation tasks to assess how H2R impacts policy learning performance in practice. For the pre-training stage, we apply MAE [he2022mae] and R3M [nair2022r3m] frameworks to train visual encoders using 62,500 videos from the SSv2 dataset, where 16 keyframes are sampled per video. The pretrained visual representations are used in downstream imitation learning pipelines. We conduct comprehensive evaluations across both simulation and real-world manipulation tasks. In simulation, we evaluate pre-trained visual encoders on four benchmark suites: Robomimic, RLBench, PushT and CortexBench. H2R improves the average success rates by +10.2% for MAE and +6.3% for R3M on Robomimic tasks, +10.0% for MAE and +5.0% for R3M on RLBench tasks, +5.3% for MAE and +7.0% for R3M on the PushT task, and +5.3% for MAE and +1.3% for R3M on the CortexBench task. In real-world experiments, we deploy H2R-enhanced encoders on different robot and end-effector settings, evaluated over nine manipulation tasks under two policy frameworks: Diffusion Policy (DP) [chi2023diffusionpolicy] and ACT [zhao2023act]. H2R brings consistent improvements, with average success rate gains of +10.0% (MAE) and +14.1% (R3M) under DP, and +12.8% (MAE) and +13.3% (R3M) under ACT. H2R also improves MAE/R3M performance when pre-trained on Ego4D, with up to +10% and +15% gains in simulated and real-world tasks, respectively. Beyond these evaluations, we also test H2R’s ability to support cross-embodiment transfer by pre-training encoders on Franka data and deploying them on UR robots, indicating that H2R improves representation transfer across different morphologies. In addition, fine-tuning VLA backbones on H2R-generated datasets yields markedly higher success rates on dual-arm UR tasks than fine-tuning on human-hand videos or using no fine-tuning. These results demonstrate that H2R effectively bridges the human-to-robot visual domain gap and significantly enhances downstream performance across both simulation and real-world manipulation settings.
2 Related Work
Robot Imitation Learning. Data-driven policy learning [sergey2018learning, luo2024hilserl, xiao2022mvp, yang2024equibotsim3equivariantdiffusionpolicy, chi2023diffusionpolicy, embodimentcollaboration2024openxembodimentroboticlearning] has enabled robots to autonomously perform tasks such as grasping, locomotion, and manipulation. Imitation learning [chi2023diffusionpolicy, ze2024dp3, zhao2023act, mandlekar2021matters] trains policies from successful demonstrations, often supervised by behavior cloning [torabi2018behavioral, florence2021implicit] objectives. ACT [zhao2023act] addresses non-Markovian dynamics by fusing temporal sequences, while diffusion models [chi2023diffusionpolicy, ze2024dp3] are introduced to handle the inherent multimodality of robot motions. In addition, CordViP [fu2025cordvip] leverages 3D object-robot correspondences to enhance dexterous manipulation.
Visual Encoder Pretraining for Robotics. Visual pretraining improves generalization of robotic policies across diverse tasks. Researchers have explored architectural designs [He_2016_CVPR, dosovitskiy12021an], training objectives [He_2022_CVPR, chen2020simple, chen2020big], and dataset compositions [deng2009imagenet, lin2014microsoft, shao2019object365, Schuhmann2021LAION400MOD]. PVR-Control [parisi22a] shows that pretrained visual representations can outperform direct state-based policies. RPT [radosavovic2023robot] tokenizes observations to enable masked prediction pretraining. Methods like MVP [radosavovic2022realworld] and R3M [nair2022r3m] utilize self-supervised objectives on videos to learn representations transferable to reinforcement learning. Voltron [karamcheti2023voltron] demonstrates the use of MAE and contrastive learning for hierarchical robot control.
Cross-Domain Visual Alignment. Bridging the domain gap between human and robot visual inputs remains a major challenge. WHIRL [bahl2022humantorobotimitationwild] matches task structure from third-person views, while RoVi-Aug [chen2024roviaugrobotviewpointaugmentation] and Mirage [chen2024mirage] manipulate appearance via segmentation or image-space preprocessing. EgoMimic [kareer2024egomimicscalingimitationlearning] removes hands and normalizes views to align egocentric perspectives.
3 H2R: Human-to-Robot Data Augmentation
In this section, we describe H2R, a data augmentation pipeline for robot learning from egocentric human videos (Figure 1). It replaces human hands in every frame with robotic arms equipped with various end effectors, generating a new, visually different dataset. This approach aims to mitigate the visual gap between human hands and robots, thereby improving the generalizability of visual representations learned from egocentric data to robotic domains. We first examine the position and pose of the hand and arm in the image, then map this pose to the robotic arm and end-effector in the simulator. Next, we erase the original hand and arm from the image and overlay a rendered image of the simulated robot performing the same pose onto the original image in place of the removed hand and arm. Figure 3 illustrates examples of human hand videos that have been processed using H2R.
3.1 H2R Data Augmentation Pipeline
Figure 1: H2R Pipeline. H2R involves replacing human hands with robotic arms by first using the HaMeR model to detect hand poses and camera parameters. The human hand is then removed using the SAM, and the inpainting model LaMa fills in the gap. A robot hand is constructed based on the detected pose and keypoints, with the camera perspective adjusted to match the original image. Finally, the robot hand is overlaid onto the image, ensuring accurate alignment with the human hand.
3D Hand Pose Estimation. In order to overlay the human hands in the egocentric image with different robots, we first need an efficient and accurate model to detect hand information. We adopt HaMeR [pavlakos2023reconstructinghands3dtransformers], a state-of-the-art model for 3D hand detection and reconstruction, to accurately locate the hand in the image, providing precise positional information for subsequent hand removal. Given an ego-centric RGB image, HaMeR estimates both the hand pose (including 3D keypoints) and the intrinsic and extrinsic parameters of the rendering camera.
Robotic Arm and End Effector Construction. This step involves building the robotic arm and its end-effectors, which are typically either grippers or dexterous hands. For dexterous hands, we estimate joint angles from hand keypoints predicted by HaMeR. Each finger joint angle is calculated using three consecutive keypoints along that finger. For grippers, we determine how open or closed they are based on the Euclidean distance between the corresponding fingertips. Since hand keypoints don’t capture the full arm pose (especially for joints unrelated to the hand), we manually set reasonable values for the remaining arm joints to complete a plausible robot configuration.
Simulator Camera Position Alignment. The visual bias introduced by the camera perspective is more significant than the action retargeting itself. To address this, we use the hand keypoints and camera parameters from HaMeR to adjust the camera pose in the simulator. Specifically, we define two coordinate systems: $C_{H}$, the coordinate system aligned with the human hand, and $C_{S}$, the coordinate system of the robotic arm in the simulator. By mapping the position of the camera in $C_{H}$ to $C_{S}$, we can ensure that the camera in the simulator shares the same perspective as the one captured in the real-world egocentric human image. The original camera position $^{W}\mathbf{cam}_{real}$ in the world frame is transformed to the aligned simulator position $^{W}\mathbf{cam}_{sim}$ using transformations from human hand ($^{W}_{H}\mathbf{R}$) and robot simulator ($^{W}_{S}\mathbf{R}$) coordinate systems:
$^{W}\mathbf{cam}_{sim}=^{W}_{S}\mathbf{R}\times^{W}_{H}\mathbf{R}^{-1}\times^{W}\mathbf{cam}_{Real}$ (1)
A more detailed explanation of the coordinate transformation and camera alignment process is provided in Appendix.
Robot Hand Rendering and Copy-paste. After setting the camera, the segmentation mask of the robotic arm is obtained by rendering with the simulator camera. We directly obtain the pixel coordinates of the human hand keypoints from HaMeR, which are predicted from the input image. In parallel, the pixel coordinates of the robot end-effector links are computed in the simulator by projecting their 3D positions through the aligned camera using the known transformation matrices. By aligning the robot link positions with the corresponding human hand keypoints in pixel space, we ensure that the overlaid robot hand accurately matches the position and orientation of the original hand in the image, achieving precise pixel-level alignment.
Human Arm and Hand Removal. After obtaining the hand pose information, we need to mask out the human hand and then inpaint the masked area with a proper background. We use Segment Anything Model (SAM) [kirillov2023segment] to automatically segment the human hand and arm regions using the hand pose information detected by HaMeR. Then, to obtain clean backgrounds for inserting robotic arms, we apply LaMa [suvorov2021resolutionrobustlargemaskinpainting], a powerful inpainting model, to fill in the removed hand-arm region. This yields clean RGB images without human limbs, providing a seamless background for inserting robotic arms in the subsequent steps.
3.2 Data Quality Evaluation
Figure 2: H2R samples. Visual comparison between original human data (top) and our augmented data (bottom).
To assess the visual plausibility and semantic consistency of the robot-augmented images produced by H2R, we employ a vision-language similarity evaluation based on CLIP [radford2021learningtransferablevisualmodels]. This method measures how well the rendered robot actions align with high-level semantic descriptions of manipulation tasks. For each augmented frame, we formulate a pair of textual prompts describing the same action: one from a human-centric perspective and one from a robot-centric perspective. The human-centric template is defined as: A human is [action]; the robot-centric template is defined as: A robotic arm is [action]. Here, [action] is a natural language phrase describing the high-level behavior, such as ”holding a bottle” or ”brushing paint on a wall”.
We use the CLIP ViT-B/32 model to compute the cosine similarity between the image embedding of the augmented frame and the corresponding textual prompt. This yields a scalar score reflecting how semantically consistent the image content is with the robot-centric description. Similarly, we compute the similarity between the original human frame and the human-centric prompt. This setup allows us to directly compare semantic alignment before and after augmentation under their respective modalities. For each frame, only the action-relevant region is retained, and background context is preserved via inpainting using LaMa. We average similarity scores over multiple samples per action class to mitigate variance and isolate the impact of robotic replacement. All textual prompts are manually curated to maintain consistency across frames and embodiments.
Figure 2 contains six pairs of subfigures, each showing a human action and its corresponding augmented robotic action. The six action pairs in Figure 2 (left to right) are described as follows: (1) Holding a yellow measuring tape at a construction site; (2) Holding a bottle; (3) Painting on the wall with a brush; (4) Rushing paint on a wall; (5) Turning on the kitchen faucet; (6) Preparing painting materials on a table.
The CLIP similarity scores are reported in Table 1. Across all samples, H2R-augmented images consistently achieve higher similarity with robot-centric action descriptions than original images do with human-centric ones. This confirms that H2R augmentation effectively improves visual-semantic alignment between human demonstrations and robotic execution.
Table 1: CLIP similarity scores. Higher values indicate better alignment between images and action descriptions.
Img-1 Img-2 Img-3 Img-4 Img-5 Img-6
Ori 30.6 23.7 31.2 31.7 27.5 28.7
Aug 32.6 28.9 32.3 32.0 28.7 31.0
This evaluation provides a lightweight, scalable, and interpretable metric for assessing how well visual hand-to-robot transformations preserve task semantics. To quantify data quality, we sample 1,000 image pairs (original and H2R-augmented). For each original image, we use Qwen-2.5-VL [bai2025qwen25vltechnicalreport] to generate a verb-noun action description (e.g., “washing bowl”). We then form two prompts: “A human is [action]” and “A robot arm is [action]”. Using CLIP ViT-B/32, we compute the cosine similarity between: (1) the human prompt and the original image, and (2) the robot prompt and the H2R-augmented image. We report the average similarity over all 1,000 samples. The raw images achieve an average score of 28.01, while H2R-augmented images reach 29.83. This demonstrates that H2R-augmented images exhibit stronger semantic alignment with robot-centric action descriptions, confirming improved visual-semantic coherence for robotic manipulation. The results validate our data curation pipeline as effective for generating high-quality, robot-relevant visual data at scale.
Figure 3: Examples of H2R Augmentation. Each column shows images before and after augmentation. The top row uses UR5 Leaphand, the middle row uses UR5 Gripper, and the bottom row uses Franka to replace human hands.
4 Simulation Experiment
Encoder Pre-training. We adopt the MAE [he2022mae, xiao2022mvp] and R3M [nair2022r3m] frameworks for pre-training, each employing a Vision Transformer (ViT) Base [dosovitskiy2021imageworth16x16words] model as the visual encoder. The training dataset used is a subset of Ego4D (117K clips) or SSv2 (1M images). Both MAE and R3M are pre-trained with $\timesA8$00 GPUs. MAE employs 800 epochs with 128 batch size and 4e-4 learning rate, while R3M uses 20K steps with 256 batch size and 1e-4 learning rate.
Policy Training. We train for 200 steps and report the mean success rate for Robomimic tasks. For tasks in RLBench [james2019rlbench], we train for 800 epochs for each policy model and test them 20 times in the RLBench environment with a random initialization. For the PushT task, we train the Diffusion Policy model for 200 epochs and report the success rate in the simulation environment. For VC-1 Cortex benchmark, we train policies using behavior cloning composed of a 3-layer MLP for 100 epochs. The training hyperparameters used in this work are identical to those described in the original paper.
Simulation Benchmark. For each pre-training method, we evaluate the performance of pre-trained encoders in imitation learning. Specifically, we select a total of seven simulation tasks in different environments, which are from Robomimic [mandlekar2021matters], RLbench [james2019rlbench], and Diffusion Policy [chi2023diffusionpolicy]. In particular, for Robomimic, we train the policies using the behavior cloning (BC) and evaluate them on tasks such as MoveCan, Square, and Lift, where the robot performs actions such as moving or lifting objects. For RLBench, we train the policies with Diffusion Policy and evaluate them on three manipulation tasks: Close Box, Close Laptop Lid and Toilet Seat Down. We use the PushT task in the Diffusion Policy evaluation framework, which evaluates a robot’s ability to push an object to a target location. We also use three MetaWorld tasks: Assembly, ButtonPress and Hammer from VC-1 Cortex benchmark for evaluation. An overview of these simulation tasks is shown in Appendix.
Table 2: Simulation Benchmark Results. Success rates (%) across imitation learning tasks in Robomimic, PushT, RLBench and CortexBench, evaluated with MAE and R3M encoders before and after applying H2R. Blue represents an increase in task success rate, while Red represents a decrease. All subsequent tables follow the same rule.
Robomimic Diffusion Policy RLBench CortexBench MoveCan Square Lift Average PushT CloseBox CloseLaptopLid ToiletSeatDown Average Assembly ButtonPress Hammer Average MAE (SSv2) 54.0 25.5 94.5 58.0 59.2 0.0 10.0 0.0 3.3 84.0 80.0 96.0 86.7 MAE (H2R-SSv2) 79.5 (+25.5%) 29.5 (+4.0%) 95.5 (+1.0%) 68.2 (+10.2%) 64.5 (+5.3%) 5.0 (+5.0%) 15.0 (+5.0%) 20.0 (+20.0%) 13.3 (+10.0%) 88.0(+4.0%) 88.0(+8.0%) 100.0 (+4.0%) 92.0(+5.3%) R3M (SSv2) 59.5 20.5 85 55 15 0 20 10 10 76.0 56.0 88 73.3 R3M (H2R-SSv2) 61.5 (+2.0%) 37.5 (+17.0%) 85.0 (0.0%) 61.3 (+6.3%) 22.0 (+7.0%) 5.0 (+5.0%) 20.0 (0.0%) 20.0 (+10.0%) 15.0 (+5.0%) 68.0 (-8.0%) 60.0 (+4.0%) 96.0 (+8.0%) 74.7 (+1.3%)
4.1 Simulation Results
Table 2 shows encoders trained on augmented data H2R-SSv2 significantly outperform those on original SSv2 across all simulation tasks. For Robomimic tasks, MAE achieves a 10.2% average success rate improvement and R3M 6.3%, with MAE seeing a substantial 25.5% gain on MoveCan. On PushT, H2R boosts MAE by 5.3% and R3M by 7.0%. In RLBench, MAE improves its performance by 10.0% on average and R3M by 5.0%. In CortexBench, H2R improves MAE by 5.3% and R3M by 1.3%. These results confirm that the H2R-augmented dataset delivers superior visual representations for imitation learning to those learned from large-scale human video datasets like SSv2, as H2R  bridges the visual gap between human hands and robotic arms to enhance vision encoder pre-training performance.
Table 3: Results on Ego4D Dataset. Success rates (%) across imitation learning tasks in the PushT and RLBench environments.
Diffusion Policy RLBench PushT CloseBox CloseLaptopLid ToiletSeatDown Average MAE (Ego4D) 51.3 0 0 5 1.7 MAE (H2R-Ego4D) 53.5 (+2.2%) 10 (+10%) 5 (+5.0%) 0 (-5.0%) 5 (+3.3%) R3M (Ego4D) 13.6 10 5 5 6.7 R3M (H2R-Ego4D) 13.6 (+0.0%) 15 (+5.0%) 5 (0.0%) 15.0 (+10.0%) 11.7 (+10.0%)
In addition to SSv2 dataset, we also perform the same experiments on the PushT task and the tasks in RLbench benchmark using the Ego4D dataset. The experimental results in the simulator are shown in Table 3, which indicates that H2R remains highly effective when applied to the Ego4D dataset.
4.2 Pretraining Model on Robotic Datasets
Table 4: Comparison with model pretrained on robotic datasets. All visual encoder weights are initialized using the R3M pretraining method, with the “R3M-H2R” variant pretrained on the H2R-augmented SSv2 dataset.
Robomimic MoveCan Square Lift Average R3M 59.5 20.5 85.0 55.0 R3M-H2R 61.5 37.5 85.0 61.3 R3M-DROID 54.0 22.0 96.0 56.7
To further evaluate how H2R compares to models pretrained directly on robotic datasets, we refer to a recent study [jiang2024robotspretrainrobotsmanipulationcentric], which conducted a systematic comparison between visual representations pretrained on large-scale egocentric human video datasets and those trained on robotic datasets such as DROID [khazatsky2024droidlargescaleinthewildrobot]. The experimental results, summarized in Table 4, show that models pretrained on robotic datasets often provide limited or even degraded performance on downstream manipulation Robomimic benchmarks. Moreover, despite being smaller and simulated rather than real-robot data, the H2R dataset outperforms the DROID dataset in pre-training effectiveness, indicating its strong potential for robotic model pre-training and reflecting its high data quality.
4.3 Effects of Demonstration Density
The relatively low success rates observed in Table 2 are primarily due to our use of keypoint frames in RLBench tasks. This approach, widely adopted in the RLBench benchmark [goyal2023rvt, jia2024lift3dfoundationpolicylifting, shridhar2023peract], accelerates both training and evaluation by reducing redundant frames. However, selecting only keyframes inevitably decreases the density of demonstration data. Since the diffusion policy employed in our imitation learning framework benefits from dense and temporally continuous data due to its long observation horizon, sparse sampling may limit its performance.
To further investigate demonstration density impact and validate H2R’s robustness, we retrain the diffusion policy with a dense RLBench dataset under identical settings. Table 5 shows higher data density significantly boosts overall success rate (e.g., MAE average from 3.3% to 45.0%). Critically, H2R consistently enhances performance across sparse and dense data regimes, confirming its general effectiveness regardless of demonstration density.
Table 5: Effects of Demonstration Density. Success rates (%) across imitation learning tasks in the RLBench environments with sparse data and dense data.
Data Model RLBench CloseBox CloseLaptopLid ToiletSeatDown Average Sparse data MAE 0.0 10.0 0.0 3.3 MAE+H2R 5.0 (+5.0%) 15.0 (+5.0%) 20.0 (+20.0%) 13.3 (+10.0%) Dense data MAE 50.0 40.0 45.0 45.0 MAE+H2R 60.0 (+10.0%) 45.0 (+5.0%) 60.0 (+15.0%) 55.0 (+10.0%)
5 Real World Experiment
Table 6: Real-world Task Results. We report the success rate (%) over real-world tasks for MAE and R3M.
Policy Task MAE (SSv2) MAE (H2R-SSv2) R3M (SSv2) R3M (H2R-SSv2) Policy Task MAE (SSv2) MAE (H2R-SSv2) R3M (SSv2) R3M (H2R-SSv2) DP Gripper-PickCube 45 65 (+20%) 40 50 (+10%) ACT Gripper-PickCube 25 30 (+5%) 25 30 (+5%) Gripper-Stack 50 55 (+5%) 55 70 (+15%) Gripper-Stack 20 35 (+15%) 20 40 (+20%) Gripper-CloseBox 55 50 (-5%) 45 65 (+20%) Gripper-CloseBox 35 40 (+5%) 40 50 (+10%) Average 50 56.7 (+6.7%) 46.7 61.7 (+15%) Average 26.7 35 (+8.3%) 28.3 40 (+11.7%) Leaphand-GraspChicken 40 55 (+15%) 10 35 (+25%) Leaphand-GraspChicken 45 50 (+5%) 10 35 (+25%) Leaphand-StandCup 35 60 (+25%) 20 50 (+30%) Leaphand-StandCup 25 50 (+25%) 20 60 (+40%) Leaphand-OpenBox 45 65 (+20%) 40 45 (+5%) Leaphand-OpenBox 30 40 (+10%) 15 20 (+5%) Average 40 60 (+20%) 23.3 43.3 (+20.0%) Average 33.3 46.7 (+13.3%) 15 38.3 (+23.3%) Franka-PlaceToy 10 15 (+5%) 10 25 (+15%) Franka-PlaceToy 25 40 (+15%) 20 30 (+10%) Franka-SweepRubbish 15 15 (0%) 20 20 (0%) Franka-SweepRubbish 20 35 (+15%) 20 20 (0%) Franka-WeightSauce 40 45 (+5%) 20 30 (+10%) Franka-WeightSauce 30 35 (+25%) 20 25 (+5%) Average 21.7 25 (+3.3%) 17.7 25 (+7.3%) Average 25 36.7 (+16.7%) 20 25 (+5%)
Table 7: Real-world Success Rates with H2R. We report the task success rates (%) using two robot embodiments (UR5, Franka) for data augmentation during pre-training. All models are evaluated on downstream tasks using the Leaphand end-effector, under MAE and R3M encoders with DP and ACT policies. Despite embodiment mismatch, H2R provides consistent gains over the baseline without augmentation.
Policy Task MAE (SSv2) MAE (H2R-UR5-SSv2) MAE (H2R-Franka-SSv2) R3M(SSv2) R3M (H2R-UR5-SSv2) R3M (H2R-Franka-SSv2) DP Leaphand-GraspChicken 40 55 (+15%) 35 (-5%) 10 35 (+25%) 20 (+10%) Leaphand-StandCup 35 60 (+25%) 50 (+15%) 20 50 (+30%) 30 (+10%) Leaphand-OpenBox 45 65 (+20%) 45 40 45 (+5%) 45 (+5%) Average 40 60 (+20%) 43.3 (+3.3%) 21.7 43.3 (+21.7%) 31.7 (+10%) ACT Leaphand-GraspChicken 45 50 (+5%) 50 (+5%) 10 35 (+25%) 40 (+30%) Leaphand-StandCup 25 50 (+25%) 50 (+25%) 20 60 (+40%) 25 (+5%) Leaphand-OpenBox 30 40 (+10%) 30 15 20 (+5%) 5 (-10%) Average 33.3 46.7 (+13.3%) 43.3 (+10%) 15 38.3 (+23.3%) 23.3 (+8.3%)
Robot Setups. We evaluate the effectiveness of H2R across three real-world manipulation setups: (i) a UR5 arm equipped with a Robotiq Gripper [RobotiqAdaptiveGrippers], (ii) a UR5 arm equipped with a Leaphand end effector [shaw2023leaphandlowcostefficient], and (iii) a dual-arm Franka Emika system using parallel Grippers. Realsense [IntelRealSenseL515] is mounted on the side of the robotic arm, which provides a similar viewpoint to the ego-centric human video data used in the pre-trained visual model. We provide the visualization of the robot setup in real-world scenarios in the supplementary materials.
Figure 4: Visualization of Real-world Manipulation Tasks. The left columns show Gripper tasks and the right columns show Leaphand tasks. Each task is illustrated with three frames, demonstrating the progression from the initial state to the completion of the manipulation.
Real-world Tasks. We set up nine tasks for gripper manipulation, dexterous manipulation and dual-arm manipulation (Figure 4). For the Gripper tasks, we design: (1) Gripper-PickCube, where the Gripper picks up a cube and places it into a bowl; (2) Gripper-Stack, where a blue cube is stacked atop a yellow cube; and (3) Gripper-CloseBox, where a cube is retrieved from a box, placed into a bowl, and the box lid is subsequently closed. For the Leaphand tasks, we design: (4) Leaphand-GraspChicken, where a toy chicken is grasped and placed into a bowl; (5) Leaphand-StandCup, where a fallen cup is stood upright on the table; and (6) Leaphand-OpenBox, where an articulated box lid is opened. For the dual-arm manipulation tasks, we design: (7) Dualarm-PlaceToy, where the left arm takes a toy from the pink box and the right arm moves it to the blue box; (8) Dualarm-SweepRubbish, where the left arm sweeps the trash into the dustpan held by the right arm; and (9) Dualarm-WeightSauce, where the left arm places the cup on the scale and the right arm picks up the bottle for a pouring motion. We adopt varying numbers of demonstrations, episode lengths per demonstration, and maximum action steps during evaluation for each task, as in Table 8.
Table 8: Task-specific Data Collection and Evaluation Settings.
Task Episode Length Num Demos Max Steps Gripper-PickCube 45 30 80 Gripper-Stack 43 30 80 Gripper-CloseBox 68 30 120 Leaphand-GraspChicken 150 50 500 Leaphand-StandCup 150 50 500 Leaphand-OpenBox 200 50 800 Dualarm-PlaceToy 1000 300 1500 Dualarm-SweepRubbish 1000 300 1500 Dualarm-WeightSauce 1000 300 1500
Table 9: Ablation Study. Ablation by removing robot overlay (w/o Overlay) and camera-hand retargeting (w/o Retarget).
Policy Task H2R H2R w/o Overlay H2R w/o Retarget DP Leaphand-GraspChicken 55 30 (-25%) 30 (-25%) Leaphand-StandCup 60 40 (-20%) 55 (-5%) Leaphand-OpenBox 65 20 (-45%) 45 (-20%) Average 60 30 (-30%) 43.3 (-16.7%) ACT Leaphand-GraspChicken 50 25 (-25%) 45 (-5%) Leaphand-StandCup 50 35 (-15%) 30 (-20%) Leaphand-OpenBox 40 25 (-15%) 15 (-25%) Average 46.7 28.3 (-18.3%) 30 (-16.7%)
Policy Training. For policy training, we select the Diffusion Policy (DP) [chi2023diffusionpolicy] and ACT [zhao2023act] as policy frameworks. We apply the pre-trained MAE and R3M visual encoders to downstream policy learning, following the same pretraining configuration described in Section 4.
Evaluation. To ensure that the initial conditions are consistent throughout the evaluation, we randomly place the target objects within a predefined area following a uniform distribution during expert demonstrations. During the evaluation, each real-world task is rolled out 20 times. We report the success rates of these tasks to assess model quality.
5.1 Real-world Results
In the real-world tasks, we employ diffusion policy (DP) [chi2023diffusionpolicy] and ACT [zhao2023act] to drive robots in performing manipulation tasks, where the visual encoders are pre-trained using MAE and R3M. To evaluate the effectiveness and generalization of H2R, we first conduct experiments to validate its performance enhancement when the pre-training and policy training embodiments are the same. We then perform cross-embodiment experiments to investigate the transfer effects of H2R when the pre-training and policy training embodiments differ. We finally apply H2R to fine-tune a pre-trained vision backbone in a Vision-Language-Action model, showcasing its promising potential in this advanced setting.
For identical embodiment experiment, we use UR5 or Franka robot for augmentation during pre-training, which matches the embodiment used in policy training and evaluation. As shown in Table 6, H2R provides consistent improvements across both MAE and R3M encoders for both DP and ACT policies. For Leaphand tasks under DP, H2R leads to an average improvement of 20% (MAE) and 21.7% (R3M). For Gripper tasks, improvements reach 6.7% (MAE) and 15% (R3M). For Dual-arm Franka tasks, the improvements reach 3.3% (MAE) and 7.3% (R3M). Similar trends are observed in ACT, with Leaphand performance increasing by 13.3% (MAE) and 23.3% (R3M), Gripper tasks improving by 8.3%(MAE) and 11.7%(R3M), and Dual-arm Franka tasks improving by 16.7%(MAE) and 5%(R3M).
To evaluate cross-embodiment generalization, we use Franka-augmented the SSv2 dataset with H2R pipeline. Downstream policy training and evaluation are conducted with UR5 Leaphand. As shown in Table 7, H2R still outperforms raw Ego4D despite the embodiment mismatch. Under DP, MAE improves from 40% to 43.3%, and R3M from 21.7% to 31.7%. Under ACT, MAE increases from 33.3% to 43.3%, and R3M from 15% to 23.3%.
5.2 Performance on Other Datasets
Table 10: Real-world Success Rates with Ego4D Pre-training. We report success rates (%) on real-world Leaphand tasks using visual encoders pre-trained on the Ego4D dataset. All downstream policies are trained using ACT.
Task MAE (Ego4D) MAE (H2R-Ego4D) R3M (Ego4D) R3M (H2R-Ego4D) Leaphand-GraspChicken 25 35 (+10%) 15 20 (+5%) Leaphand-StandCup 25 50 (+25%) 20 25 (+5%) Leaphand-OpenBox 35 45 (+10%) 30 40 (+10%) Average 28.3 43.3 (+15%) 21.7 28.3 (+6.7%)
We compare encoders pre-trained on the original Ego4D dataset and on the H2R-augmented dataset, and apply them to real-world Leaphand manipulation tasks using the ACT [zhao2023act] policy. As shown in Table 10, models trained with H2R data consistently outperform those using raw Ego4D across all three Leaphand manipulation tasks, achieving an average improvement of 15% for MAE and 6.7% for R3M. These results complement our findings on SSv2 and reinforce that robot-centric augmentation improves downstream real-world performance.
5.3 Potential for VLA Model Pretraining
Figure 5: Dual-UR5 tasks. The left columns shows PlaceBowl task and the right columns shows StackBlocks task.
A common paradigm in Vision-Language-Action (VLA) models is to initialize the visual backbone using encoders pre-trained on large-scale datasets (e.g., CLIP, VAEs). In this experiment, we investigate whether fine-tuning a pre-trained visual backbone on robot-oriented video data can enhance its suitability for robotic policy learning. Specifically, we fine-tune the pre-trained VAE encoder and decoder used in the Unified Video Action Model(UVA) [li2025unifiedvideoactionmodel] separately on both the original SSv2 and the H2R-augmented datasets. The fine-tuned VAEs are subsequently frozen and integrated into the UVA framework for policy training on two distinct dual-arm UR5 robot tasks(shown in Figure  5). The results shown in Table  11 demonstrate a clear performance trend in the different variants of the UVA model.
Table 11: Success rates of UVA variants with different visual backbones.
Model PlaceBowl StackBlocks Average UVA 0.2 0.2 0.2 UVA + SSv2 0.1 0.2 0.15 UVA + SSv2 + H2R 0.4 0.3 0.35
5.4 Ablation Study
To evaluate the effectiveness of each component in H2R, we conduct ablation studies on two time-consuming steps: (1) performing hand inpainting without overlaying a robotic arm (H2R w/o Overlay), and (2) overlaying the arm without precise alignment between the hand and the camera, instead using random pasting (H2R w/o Retarget). Table 9 shows the necessity and effectiveness of each component in H2R. The first step leads to a significant drop in success rate due to the loss of critical human-object interaction pixels after inpainting. The second step fails to provide accurate motion cues for the model and introduces visual mismatches with real-world manipulation tasks.
6 Conclusion
We propose H2R, a data augmentation technique that bridges the visual gap between human hand demonstrations and robotic arm manipulations by replacing human hands in first-person videos with robotic arm movements. Using 3D hand reconstruction and image inpainting models, H2R generates synthetic robotic arm manipulation sequences, making them more suitable for robot pre-training. Experiments across simulation benchmarks and real-world tasks demonstrate consistent improvements in success rates for encoders trained with various pre-training methods (e.g., MAE, R3M), highlighting the effectiveness and generalization of H2R. H2R enables efficient transfer of task knowledge from human demonstrations to robotic systems, reducing the reliance on costly robot-specific data collection.
[…]
Supplementary Material
Appendix A Details of Simulator Camera Position Alignment
We define two coordinate systems: $C_{H}$, the coordinate system aligned with the human hand, and $C_{S}$, the coordinate system of the robot arm in the simulator. We build the coordinate system $^{W}\mathbf{I}_{H}$ based on the hand keypoints:
$^{W}\mathbf{I}_{H}=\{^{w}\mathbf{i}_{H,x},^{w}\mathbf{i}_{H,y},^{w}\mathbf{i}_{H,z}\}$ (2)
Where $^{w}\mathbf{i}_{H,x},^{w}\mathbf{i}_{H,y},^{w}\mathbf{i}_{H,z}$ are unit vectors along the x-axes, y-axes and z-axes of the human hand coorinate system. With the keypoints get in HaMeR, we build the three axis of coordinates with the following functions:
${}^{w}\mathbf{i}_{H,x}$ $=^{w}\mathbf{i}_{0,9}$ (3)
${}^{w}\mathbf{i}_{H,y}$ $=^{w}\mathbf{i}_{0,9}\times^{w}\mathbf{i}_{0,13}$
${}^{w}\mathbf{i}_{H,z}$ $=^{w}\mathbf{i}_{H,x}\times^{w}\mathbf{i}_{H,y}$
Where $^{w}\mathbf{i}_{0,9}$ and $^{w}\mathbf{i}_{0,13}$ are unit vectors along the middle and ring fingers, respectively. In this notation, the first index (0) refers to the specific finger (middle or ring), and the second index (9 and 13) corresponds to the joint numbers along those fingers, as defined by the MANO model. Similarly, To construct the mapping from hand pose to robot arms, we need to get another coordinate system $^{W}\mathbf{I}_{S}$ in the simulator:
$^{W}\mathbf{I}_{S}=\{^{w}\mathbf{i}_{S,x},^{w}\mathbf{i}_{S,y},^{w}\mathbf{i}_{S,z}\}$ (4)
The method of determining the axis of coordinates is the same:
${}^{w}\mathbf{i}_{S,x}$ $=^{w}\mathbf{i}_{0,2}$ (5)
${}^{w}\mathbf{i}_{S,y}$ $=^{w}\mathbf{i}_{0,2}\times^{w}\mathbf{i}_{0,3}$
${}^{w}\mathbf{i}_{s,z}$ $=^{w}\mathbf{i}_{S,x}\times^{w}\mathbf{i}_{S,y}$
Where $\mathbf{i}_{0,2},\mathbf{i}_{0,3}$ are unit vectors along robot fingers that correspond to human middle and ring fingers and the index corresponds to the joint numbers defined by MANO. We build the following two coordinate transformation matrix to construct the mapping:
${}^{W}_{H}\mathbf{R}$ $=\begin{pmatrix}{}^{W}\mathbf{I}_{H}&\mathbf{key}_{human}\\ \mathbf{O}&1\end{pmatrix}$ (6)
${}^{W}_{S}\mathbf{R}$ $=\begin{pmatrix}{{}^{W}\mathbf{I}_{S}}&{\mathbf{key}_{robot}}\\ \mathbf{O}&{1}\end{pmatrix}$
Where $\mathbf{key}_{human},\mathbf{key_{robot}}$ are the positions of human wrist and robot wrist. After obtaining the two coordinate systems, we need to determine the position of the camera in the simulator ($^{W}\mathbf{cam}_{sim}$) and the position of the camera in the real world ($^{H}\mathbf{cam}_{Real}$), thus we can ensure we get the same pose of the human hand and robot arms
${}^{H}\mathbf{cam}_{Real}$ $=^{W}_{H}\mathbf{R}^{-1}\times^{W}\mathbf{cam}_{Real}$ (7)
${}^{S}\mathbf{cam}_{sim}$ $=^{H}\mathbf{cam}_{Real}$
${}^{W}\mathbf{cam}_{sim}$ $=^{W}_{S}\mathbf{R}\times^{W}_{H}\mathbf{R}^{-1}\times^{W}\mathbf{cam}_{Real}$
Appendix B Details of Experimental Setup
Figure 6: Simulation benchmark. We choose 3 tasks from the Robomimic, 3 tasks from the RLBench, and 3 tasks from the CortexBench, covering a range of robotic manipulation skills. We also include the PushT task, designed for the Diffusion Policy framework, as an additional benchmark to evaluate performance in a different task setup.
Figure 7: Real-world Robot Setup. Illustration of different real-world experimental setup.
Data Collection in Real-world Experiment. For the UR5-gripper, demonstration trajectories were collected via keyboard-based teleoperation. For the UR5-Leaphand, demonstrations were gathered using vision-guided teleoperation. For the Dual-arm Franka and Dual-arm UR5 setups, data collection was performed through homogeneous-arm bilateral teleoperation.
Additional Evaluation Details. Figure  6 provides visualizations of the different simulation benchmarks, and Figure  7 illustrates the real-world experimental setup. For the simulation benchmarks, we use the default configurations for both training and evaluation. In the real-world tasks, the UR5–Gripper and UR5–Leaphand setups use the camera shown in Figure  7 as the visual input, whereas the dual-arm Franka and dual-arm UR5 configurations use a front–top camera together with two wrist-mounted cameras (three cameras in total) for visual observation. To ensure that the initial conditions are evaluated consistently across algorithms, we randomly place the target objects within a predefined area following a uniform distribution during expert demonstration collection. During evaluation, the object placement region is kept identical to that used in data collection, ensuring a fair and consistent comparison.
Stack
Stand Cup
Figure 8: Failure case visualizations: Stack and Stand Cup. We visualize real-world manipulation executions for two downstream tasks: Stack (top) and Stand Cup (bottom). These images provide qualitative insights into the performance and failure modes of the policy in real deployment, highlighting challenges such as object misalignment, perception noise, and grasp precision.
Appendix C Failure Case Analysis
Table 12: Sub-goal.Task-specific sub-goal evaluation. To gain fine-grained insights into policy performance, we design a manual rubric covering key sub-goals for each manipulation task. Each cell reports the number of successful vs. unsuccessful attempts (Y/N) over 20 evaluation trials. Results show that models enhanced with H2R consistently accomplish more sub-goals across tasks compared to their baseline counterparts, demonstrating improved robustness in real-world execution. Bold numbers indicate better performance between paired models.
Task Sub-goal MAE(Y/N) MAE+H2R(Y/N) R3M(Y/N) R3M+H2R(Y/N) Gripper-PickCube Overall success? 9/11 13/7 8/12 10/10 Pick up the cube? 14/6 15/5 11/9 13/7 Gripper-Stack Overall success? 10/10 11/9 11/9 14/6 Pick up the cube? 13/7 16/4 13/7 17/3 Gripper-CloseBox Overall success? 11/9 10/10 9/11 13/7 Place the cube in the bow? 12/8 14/6 12/8 15/5 Pick up the cube? 14/6 14/6 12/8 15/5 Leaphand-GraspChicken Overall success? 8/12 11/9 2/18 7/13 Pick up the chicken? 13/7 14/6 3/17 10/10 Leaphand-StandCup Overall success? 7/13 12/8 4/16 10/10 Pick up the cup? 12/8 18/2 12/8 15/5 Leaphand-OpenBox Overall success? 9/11 13/7 8/12 9/11 Identify contact location? 14/6 16/4 10/10 10/10
To better understand the limitations of our policy and the challenges encountered in real-world deployments, we present a qualitative analysis of failure cases from two representative tasks: a gripper-based task (Gripper-Stack) and a dexterous manipulation task (Gripper-StandCup). Figure 8 illustrates typical failure modes observed during execution.
In the Gripper-Stack task, we identify two major failure scenarios:
Case I: Grasp Failure Unnoticed. The robot arm fails to successfully grasp the blue cube. However, the policy proceeds as if the object had been grasped, moving toward the yellow cube and attempting to perform the stacking operation. This leads to a complete task failure.
Case II: Misaligned Placement. The robot successfully grasps the blue cube but fails to align it correctly on top of the yellow cube during the stacking phase, resulting in an unstable or failed placement.
In the Stand Cup task, similar issues emerge due to perception and control limitations:
Case I: Grasp Position Error. The Leaphand end-effector attempts to grasp the cup but fails to target the correct contact region. As a result, the cup slips out of the grasp during lifting, preventing task completion.
Case II: Insufficient Lifting Trajectory. Even when the grasp is successful, the lifting motion lacks sufficient amplitude or stability to fully stand the cup upright. The cup either tips over or fails to stand securely.
To enable fine-grained evaluation of policy performance and gain deeper insights into failure cases, we designed a task-specific evaluation rubric. Table 12 displays our rubric that the evaluator filled out when rolling out different policies. Take the DP policy as an example, the results in Table 12 demonstrate that H2R-augmented visual representation models not only improve overall success rates in real-world tasks, but also allow to accomplish more than half of the task consistently.
Table 13: Robomimic Experiment result. We report the success rate (%) over IL-based tasks for MAE and R3M Robomimic.
MoveCan Square Lift Average PushT MAE 54 25.5 94.5 58 59.2 MAE+CutMix1 72.0 (+18.0%) 30.0 (+4.5%) 95.0 (+0.5%) 65.7 (+7.7%) 37.5 (-21.7%) MAE+CutMix2 58.0 (+4.0%) 36.0 (+10.5%) 90.0 (-4.5%) 61.3 (+3.3%) 40.0 (-19.2%) MAE+CutMix3 78.0 (+24.0%) 32.0 (+9.3%) 92.0 (-2.5%) 67.3 (+2.7%) 42.0 (-17.2%) MAE+H2R 79.5 (+25.5%) 29.5 (+4.0%) 95.5 (+1.0%) 68.2 (+10.2%) 64.5 (+5.3%) R3M 59.5 20.5 85 55 15 R3M+CutMix1 69.5 (+10.0%) 30.0 (+9.5%) 91.0 (+6.0%) 63.5 (+8.5%) 19.0 (+4.0%) R3M+CutMix2 66.0 (+6.5%) 26.0 (+5.5%) 83.0 (-2.0%) 58.3 (+3.3%) 17.0 (+2.0%) R3M+CutMix3 68.0 (+8.5%) 26.0 (+5.5%) 84.0 (-1.0%) 59.3 (+4.3%) 14.0 (-1.0%) R3M+H2R 61.5 (+2.0%) 37.5 (+17.0%) 85.0 (0.0%) 61.3 (+6.3%) 22.0 (+7.0%)
Table 14: Generalization under Lighting Variations. Success rates (%) under real-world lighting disturbances. Models trained with additional lighting augmentations (LightAug) show significant improvements in robustness compared to baseline and H2R-only models.
Tasks MAE MAE+H2R MAE+H2R+LightAug R3M R3M+H2R R3M+H2R+LightAug Leaphand-GraspChicken 10 10 20 0 0 10 Leaphand-StandCup 20 25 40 5 15 20 Leaphand-OpenBox 5 0 25 10 10 10 Average 11.7 11.7 28.3 5 8.3 13.3
Appendix D Additional Ablation Study in Simulation Experiment
In addition to pre-training on the H2R data and raw data, we also applied a simple CutMix baseline to demonstrate the effectiveness of using the robotic arm to cover the human hand, which overlays a fixed set of specific images of robotic arms with grippers onto the original images, ensuring that the overlaid images cover the human hands as much as possible, without exceeding the detected bounding box. Our H2R is different from such baseline by employing robot hand construction to better match the pose of the hand and arm in the images. Based on the type of robotic arm used in CutMix, we categorize the augmented set into three types: CutMix1 represents the UR5 robotic arm, CutMix2 refers to the Franka robotic arm, and CutMix3 combines both the UR5 and Franka robotic arms.
From Table 13, we observe that the encoder trained on H2R processed data shows consistent improvements across various tasks compared to the encoder trained on the original data, with the average success rate on all tasks ranging from 0.9% to 10.2%. Especially for the more challenging MoveCan task, it can improve the success rate by 25.5%. Additionally, while encoders trained on the relatively simple CutMix data show improvement on tasks in Robomimic, their performance in the PushT task remains slightly worse than the encoders trained on original data. These results demonstrate the effectiveness of using the robotic arm to cover the human hand in video data, as well as the effectiveness of H2R in imitation learning.
Appendix E Generalization on Light Condition.
Figure 9: Lighting Setup in Real-World Experiments. Light configuration used in real-world evaluations.
To evaluate the generalization under varying lighting conditions, we introduce illumination disturbances during evaluation, as illustrated in Figure 9. Additionally, during training, we incorporate randomized lighting with varying directions and colors into the simulation environment for data augmentation. We compare three settings: no augmentation(MAE, R3M), H2R augmentation(MAE+H2R, R3M+H2R), and H2R with lighting disturbances(MAE+H2R+LightAug, R3M+H2R+LightAug). As shown in Table 14, the model trained with H2R and lighting perturbations demonstrates significantly better generalization to real-world lighting variations than other baselines, highlighting the effectiveness of H2R in bridging the domain gap caused by lighting variations.